OpenStudy (anonymous):

Design and implement a Java class named Sum, to produce a progression, where each value is the sum of all the previous values.

OpenStudy (anonymous):

If I understand correctly, you'll never get any result 'cause for example when the first value is "0", all the other values (= sum of previous values) would also be "0". This plays for every start value you choose. Or do you want to sum up like the following: value of "3" = 0+1+2 = 3; value of "5" = 0+1+2+3+4 = 10 ? please give an example

OpenStudy (anonymous):

Nevermind I got it! But thank you so much for the answer.